VEGAS / DIBAS Weekly Review Meeting - 2013 September 18

Connection Details:

Green Bank GB-Basement Video: 192.33.117.12##7145 Audio: 434-817-6459

Present

  • Richard, Marty,Karen, Lory

DIBAS Status

  • RMP informed SHAO of the specification change for coherent search mode - no response as yet.
  • Ray switched to DIBAS yesterday, he updated his documentation (8 hours). He's going to start picking items off the TODO list today.
  • Joe spent a few hours this week porting recent GPU fixes to the GUPPI part of DIBAS.

VEGAS Status

  • Engineering tests of mode 4 last week were unsuccessful. The results imply the BOF had not been (or had not been fully) tested.
  • On the upside, good progress was made in fixing the problems with the assistance of Glenn Jones. Engineering tests continue, and more astronomical test time is available if needed this evening.

SHAO communications status

* SHAO have asked what spectral line modes will be available for testing. RMP replied with the currently tested modes; no reply as yet.

-- RichardPrestage - 2013-09-10
Topic revision: r4 - 2013-09-18, RichardPrestage
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback